OpenStudy (anonymous):

When a mole of lithium reacts chemically with hydrogen to form lithium hydride, the energy released is found to be roughly 90 joules. How much mass is lost? 1.0 x 10^-15 kg 2.7 x 10^10 kg 1.0 x 10^15 kg 3.0 x 10^-7 kg

OpenStudy (anonymous):

E=mc^2 --> m= E/c^2 = 90 J / (3*10^8)^2

OpenStudy (anonymous):

3.0 x 10^-7 kg

OpenStudy (anonymous):

i got 1.0 x 10^-15 kg